OpenStudy (anonymous):

A pitcher throws a baseball straight into the air with a velocity of 72 feet/sec. If acceleration due to gravity is −32 ft/sec2, how many seconds after it leaves the pitcher's hand will it take the ball to reach its highest point? Assume the position at time t = 0 is 0 feet.

OpenStudy (welshfella):

use an equation for constant acceleration u = initial velocity = 72 , a = -32, t = ? , v = final velocity = 0 v = u + at 0 = 72 - 32t solve for t

OpenStudy (anonymous):

2.25?

OpenStudy (welshfella):

72 / 32 = 2.25 yes

OpenStudy (anonymous):

thank you!

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

i'm meeting so many people who dont say thanks today....

OpenStudy (anonymous):

Its common courtesy to say thank you :p

OpenStudy (anonymous):

especially for an answer

OpenStudy (welshfella):

yes